# Accelerating AllReduce with a Persistent Straggler

### v1

@@ Appendix B Proof of Communication Complexity
First, we prove that schedules synthesized by StragglAR for power-of-2 $n$ are guaranteed to complete within $n{+}\log n{-}2$ rounds. As in Algorithm 1, $A$ represents the data structure that maps each active chunk to the set of non-straggler ranks that holds it (and is updated in each round). We assume this data structure is a hash table for constant update and access. Note that an active chunk necessarily implies that the chunk has been fully reduced. Our code generates results that match the proof bounds.
@@
($n{-}2$ appears due to the straggler pairing in round $r$). Substituting the closed form of the sum,
@@
This means all $n{-}1$ non-straggler ranks have a single fully reduced chunk after Phase 1. ∎
@@
The first part of this statement follows directly from the schedule, as it dictates that partially reduced chunk $c_{r}$ resides on rank $r$ given the ReduceScatter pre-condition, and rank $r$ communicates with the straggler, $\sigma$, in round $r$, enabling $c_{r}$ to be fully reduced on ranks $r$ and $\sigma$.
@@
Base case. We rely on Lemma 1 for the base case. Immediately before round $\log n$ (after round $\log n{-}1$), $|A[c_{j}]|=2^{\log n-1-j}$, where sets $A[c_{j}]$ are pairwise disjoint for all $j=0,1,\dots,\log{n}{-}1$. Thus, $|P_{\log n}|=|A_{c_{0}}|=2^{\log n-1}=\tfrac{n}{2}$ and $|Q_{\log n}|=n{-}1{-}\tfrac{n}{2}=\tfrac{n}{2}{-}1$. Since rank $\log n$ holds chunk $c_{0}$, as the protocol in Algorithm 1 dictates that rank 00 sends $c_{0}$ to rank $\log n$ in round $1$, we know $\log n\in P_{\log n}$. By matching $u\in P_{\log n}{\setminus}\{\log n\}$ with $v\in Q_{\log n}$ and having $u$ and $v$ each send their active chunks, we ensure that $c_{0}$ has fully propagated after round $r={\log n}$. A matching is possible because every rank in $P_{\log n}$ lacks the active chunk of every rank in $Q_{\log n}$, and vice versa. After the round, $c_{0}$ has fully propagated and every other active chunk has doubled.
@@
We verify the validity of $\mathcal{I}(r{+}1)$ before round $r{+}1$ as follows. We can remove $c_{r-\log n}$ from $A$ because it has fully propagated via the matching between $P_{r}$ and $Q_{r}$ in round $r$. We add $c_{r}$ to $A$ such that $A[c_{r}]=\{r\}$, since it has just been fully reduced in round $r$ via the straggler pairing. Every other active chunk has doubled in propagation (i.e., the number of ranks that possess it) via the perfect matching between $P_{r}$ and $Q_{r}$. Let $j^{\prime}=r{+}1{-}\log n$. Immediately before round $r{+}1$ we verify:
@@
    Since each rank still holds exactly one active chunk, the sets $A[c_{k}]$ remain pairwise disjoint. In round $r$, ranks in $Q_{r}$ received $c_{r-\log n}$, which is no longer active by round $r{+}1$, so these ranks still retain the same active chunk they had prior to round $r$. Ranks in $P_{r}\setminus\{r\}$ had only active chunk $c_{r-\log n}$ before round $r$, which expired as an active chunk, and received some other active chunk in round $r$, so they also only have one active chunk each. Finally, $r\in P_{r}$, so its active chunk expired after round $r$, and it received active chunk $c_{r}$ via the straggler pairing, so it also only has one active chunk. This ensures pairwise disjointness.
There are two remaining components of the proof for the inductive step: (1) that $r{+}1\in P_{r+1}$, and (2) that there always exists a feasible matching between $P_{r+1}\setminus\{r+1\}$ and $Q_{r+1}$. First, it is guaranteed that $r{+}1\in P_{r+1}$ because of the rule for matching ranks in the critical window from Algorithm 1. Specifically, we ensure that rank $r{+}1$ is never paired with a rank whose active chunk is $c_{l}$ for $l>r{+}1{-}\log n$. This rule ensures that $r{+}1\in P_{r+1}$ because the only active chunk it could have at the beginning of round $r{+}1$ is then $c_{r+1-\log n}$, guaranteeing that $r{+}1\in P_{r+1}$.
The next question is whether we can ensure that there exists a matching in round $r{+}1$. First, we need to ensure that any rank $l$ in the critical window is not matched with a rank whose active chunk will continue to be active by round $l$. If $l\in Q_{r+1}$, this is trivial because the rank does not receive any new active chunk in that round (since it only receives the chunk that is due in that round). If $l\in P_{r+1}$, we just have to ensure that $l$ is not matched with any of the “newest” active chunks; in fact, matching $l$ with ranks possessing the second-oldest active chunk always works to keep $l$ in future $P_{r}$. Since $G_{r}$, the graph that connects $P_{r}$ and $Q_{r}$, is a complete bipartite graph, we can easily guarantee this pairing for any $l$. By removing these $l$ and their partners from their respect sets (either $P_{r}$ or $Q_{r}$), we can trivially apply Hall’s Marriage Theorem since $|P_{r}|=|Q_{r}|$ (and $G_{r}$ is bipartite) to verify that a matching exists. Hence, $\mathcal{I}(r{+}1)$ holds, completing the inductive step. ∎
@@
After round $n{-}2$, the straggler, $\sigma$, now has all fully reduced chunks, and is no longer constrained by whom to pair with. This means we can add the straggler to $Q_{r}$ for subsequent rounds $r\geq n{-}1$ such that $|P_{r}|=|Q_{r}|=\tfrac{n}{2}$. (We enforce the invariant for subsequent rounds that the straggler can only send $c_{n-2}$.) Both ranks $n{-}2$ and $\sigma$ are in $Q_{r}$ (and will remain in $Q_{r}$ until the end of the algorithm), and their active chunk is $c_{n-2}$. Thus, $c_{r}$ begins with 2 ranks holding it, as opposed to 1, jumpstarting its doubling process. Every chunk can continue to double because $|P_{r}|=|Q_{r}|$ and every rank only has one active chunk, so via Hall’s Marriage Theorem a perfect matching exists. Therefore, only $\log n-1$ rounds are required to fully propagate the final chunk.
@@
Naturally, $n{-}1$ rounds are required for all chunks to become fully reduced, via the linear sequence of pairing with the straggler, with chunk $c_{j}$ fully reduced in round $j$. Lemma 2 proves that each chunk $c_{j}$ propagates by round $j{+}\log n$. Thus, the total number of rounds in the algorithm is bounded by the final chunks. Chunk $c_{n-3}$ will propagate by round $n{-}3{+}\log n$ and chunk $c_{n-2}$ will have fully propagated after round $n{-}2{+}\log n{-}1=n{-}3{+}\log n$. Since we zero-index rounds, this results in $n{+}\log n{-}2$ rounds in total. ∎
Appendix C StragglAR for Non-Power-of-$2$ Values of $n$
@@
When the number of GPUs is not a power of $2$, StragglAR becomes more challenging to implement because GPUs can possess multiple active chunks at any point in time. To handle the even, non-power-of-$2$ case, we use a similar matching-based approach to construct the schedule in every round, but instead connect vertices $u$ and $v$ with weight $2$ if $u$ possesses a chunk that $v$ needs and $v$ possesses a chunk that $u$ needs, and with weight $1$ if only $u$ possesses a chunk that $v$ needs, but the opposite is not true (or vice versa). Then, we run well-known polynomial-time algorithms  [82, 83] to compute the maximum weight matchings per round in order to generate the schedule. Thus, offline schedule generation when $n$ is not a power-of-2 has a higher runtime, but is still polytime computable.
@@
For even, non-power-of-2 values of $n$, we do not prove a bound on the $\alpha{-}\beta$ cost. However, we still synthesize and evaluate these schedules in §4 and typically find schedules that complete in ${\sim}n{+}2\log n{-}2$ rounds in practice, which still results in lower $\beta$ cost than baselines. The realized communication complexity of the schedules generated by StragglAR for non-power-of-2 $n$ is shown in Fig. 6(b) (right), and still outperforms all baselines using the analytical network simulator. We note that StragglAR does not provide schedules for odd values of $n$, but these settings are less common in large-scale distributed ML.
Appendix D Evaluation on Perlmutter with 4 GPUs
@@
We show the results from running on one node of the Perlmutter supercomputer [16] in Figure 7. While StragglAR still consistently outperforms baselines across buffer sizes in the 4-GPU setup, its gains are less substantial than for the 8-GPU setup. This is expected, and the results are in line with the theoretical results in §4 because StragglAR’s performance advantage improves as $n$, the size of the GPU cluster, scales. (Intuitively, this is because the proportion of “work” that is offloaded to the pre-condition is higher as $n$ scales.) Figure 7b indicates similar performance on the 4-GPU node even with both the average and masked delays, suggesting the ReduceScatter precondition can be effectively masked in typical scenarios.
Appendix E NCCL Internal Chunking Behavior
As we discuss in §4, the performance of ncclSend()/ncclRecv() deviates from the $\alpha{-}\beta$ cost model’s expected performance for some specific buffer sizes due to heuristics and chunking behavior implemented internally by NCCL. We run the baseline nccl-tests [31], a benchmarking suite provided by Nvidia for evaluating in-built NCCL performance, across different data sizes and plot the results in Fig. 8. The region from 64 MiB-256 MiB deviates from expected behavior under $\alpha{-}\beta$ costs because sending 2×2$\times$2 ×, and even 4×4$\times$4 ×, the amount of data results in almost no change in communication time with the P2P API. (Under the $\alpha{-}\beta$ model, the time should consistently scale linearly with the send/recv data size, instead of step-wise as shown in Fig. 8.) This is why the Direct AllReduce baseline outperforms other algorithms at 256 MiB, as the time to send an entire 256 MiB buffer is only marginally higher than sending smaller chunk sizes, as bandwidth-efficient algorithms like Ring, RHD, and StragglAR do. We do not observe this finding to generalize to other architectures, and may be a relic of the NCCL version, CUDA version, and specific environment configured by RunPod for ncclSend()/ncclRecv(). Also, outside of the 64-256 MiB range, the performance closely reflects the $\alpha{-}\beta$ model, and our findings in §4 reflect this.
@@ Appendix E NCCL Internal Chunking Behavior
Appendix F End-to-End Training Experiments on Multi-GPU Servers
@@
We evaluate StragglAR using data-parallel (DP) training of the Llama3.2-3B model in PyTorch on the RunPod server, which consists of 8 A100 GPUs connected via NVSwitch. In this evaluation, we enable AllReduce algorithms to run in the wild, calling implementations directly from PyTorch on a real hardware testbed. To do so, we compile our C++ and CUDA implementations (using the NCCL P2P API, exactly as described in §4) into a custom package, and enable dist.all_reduce() to call these implementations using PyTorch distributed’s cpp extension capabilities [84]. We compare StragglAR to the Ring algorithm, the most well-known and widely used bandwidth-efficient baseline, under this scheme. Following the profiling methodology in §2, we identify device (GPU) 6 to be the reliable straggler on our server. At the start of training (and only once for the run), we remap this straggling device to rank 7 in the CUDA runtime to ensure StragglAR executes correctly. Then, training simply proceeds as per usual with PyTorch.
Because this is data-parallel training with a large model, there are multiple AllReduce calls that proceed sequentially at the end of each backward pass, each for a different subset of the model parameters. The first, and by far the largest (more than 30×30$\times$30 × the size of any other subset of parameters), is the embedding table. This is also the primary AllReduce that could incur a straggler because subsequent AllReduce calls benefit from the inherent synchronization resulting from the previous AllReduce with no compute blocks in between. We measure the AllReduce time for this first (bulk) AllReduce on device 0 (exactly as in §4, but this time in the PyTorch layer) for 50 training iterations in two different runs with global batch sizes of 32 and 64 each. (These batch sizes are based on typical fine-tuning use cases that opt for smaller batch sizes [85, 86, 87].) We show the average, with error bars for standard deviation in Fig. 9(a). The smaller batch size of 32 results in only 1.03×1.03$\times$1.03 × speedup due to the limited straggler delay by under-utilizing GPU memory and compute. Meanwhile, the larger batch size of 64 results in 1.13×1.13$\times$1.13 × speedup over Ring AllReduce since it maximizes GPU utilization and thereby introduces a more severe straggler effect that can overlap the ReduceScatter pre-condition.
@@
Then, we finetune Llama3.2-3B using the same data-parallel setup on 8 GPUs with the global batch size of 32 for 100 iterations to determine the overall training efficiency gains with StragglAR. Fig. 9(b) shows that even just for these first 100 iterations, StragglAR reduces training time by 7 seconds. These gains suggest that training at longer time scales—for example, weeks—could shave off almost a day of training simply by using StragglAR for AllReduce. Further, we note that the StragglAR curve is just the Ring AllReduce curve shifted to the left: the $y$-coordinates are exactly the same because both algorithms ensure complete data buffers are reduced, ensuring no change in model convergence behavior. Instead, StragglAR enables an identical, full-parameter AllReduce, but just in less time per iteration, resulting in greater training efficiency and faster time-to-convergence. Fig 9(b) captures the wall-clock time, i.e., all compute and communication operations in training, so the results may improve in more distributed settings (as shown in simulations in §G) or with even larger model sizes where communication becomes even more of a bottleneck in distributed ML.
(a) Runtime of DP AllReduce call.
(b) Training efficiency, 100 iters. ($B{=}32$).
Figure 9: End-to-end evaluation of AllReduce implementations on Llama3.2-3B finetuning with PyTorch on a RunPod server with 8 A100 GPUs connected via NVSwitch.
Appendix G End-to-End Training Simulations at Scale

### v2

@@ Appendix C Proof of Communication Complexity
First, we prove that schedules synthesized by StragglAR for power-of-2 $n$ are guaranteed to complete within $n{+}\log n{-}2$ rounds. As in Algorithm 1, $A$ represents the data structure that maps each active chunk to the set of non-straggler ranks that holds it (and is updated in each round). We assume this data structure is a hash table for constant-time updates and access. Note that an active chunk necessarily implies that the chunk has been fully reduced. Our code generates results that match the proof bounds.
@@
(n2n{-}2 appears due to the straggler pairing in round $r$). Substituting the closed form of the sum,
@@
This means all n1n{-}1 non-straggler ranks have a single fully reduced chunk after Phase 1. ∎
@@
The first part of this statement follows directly from the schedule, as it dictates that partially reduced chunk $c_{r}$ resides on rank $r$ given the ReduceScatter precondition, and rank $r$ communicates with the straggler, $\sigma$, in round $r$, enabling $c_{r}$ to be fully reduced on ranks $r$ and $\sigma$.
@@
Base case. We rely on Lemma 1 for the base case. Immediately before round $\log n$ (after round $\log n{-}1$), $|A[c_{j}]|=2^{\log n-1-j}$, where sets $A[c_{j}]$ are pairwise disjoint for all $j=0,1,\dots,\log{n}{-}1$. Thus, $|P_{\log n}|=|A_{c_{0}}|=2^{\log n-1}=\tfrac{n}{2}$ and $|Q_{\log n}|=n{-}1{-}\tfrac{n}{2}=\tfrac{n}{2}{-}1$. Since rank $\log n$ holds chunk $c_{0}$, as the protocol in Algorithm 1 dictates that rank 0 sends $c_{0}$ to rank $\log n$ in round $1$, we know $\log n\in P_{\log n}$. By matching $u\in P_{\log n}{\setminus}\{\log n\}$ with $v\in Q_{\log n}$ and having $u$ and $v$ each send their active chunks, we ensure that $c_{0}$ has fully propagated after round $r={\log n}$. A matching is possible because every rank in $P_{\log n}$ lacks the active chunk of every rank in $Q_{\log n}$, and vice versa. After the round, $c_{0}$ has fully propagated and every other active chunk has doubled.
@@
We verify the validity of $\mathcal{I}(r{+}1)$ before round r+1r{+}1 as follows. We can remove $c_{r-\log n}$ from $A$ because it has fully propagated via the matching between $P_{r}$ and $Q_{r}$ in round $r$. We add $c_{r}$ to $A$ such that $A[c_{r}]=\{r\}$, since it has just been fully reduced in round $r$ via the straggler pairing. Every other active chunk has doubled in propagation (i.e., the number of ranks that possess it) via the perfect matching between $P_{r}$ and $Q_{r}$. Let $j^{\prime}=r{+}1{-}\log n$. Immediately before round r+1r{+}1 we verify:
@@
    Since each rank still holds exactly one active chunk, the sets $A[c_{k}]$ remain pairwise disjoint. In round $r$, ranks in $Q_{r}$ received $c_{r-\log n}$, which is no longer active by round r+1r{+}1, so these ranks still retain the same active chunk they had prior to round $r$. Ranks in $P_{r}\setminus\{r\}$ had only active chunk $c_{r-\log n}$ before round $r$, which expired as an active chunk, and received some other active chunk in round $r$, so they also only have one active chunk each. Finally, $r\in P_{r}$, so its active chunk expired after round $r$, and it received active chunk $c_{r}$ via the straggler pairing, so it also only has one active chunk. This ensures pairwise disjointness.
There are two remaining components of the proof for the inductive step: (1) that $r{+}1\in P_{r+1}$, and (2) that there always exists a feasible matching between $P_{r+1}\setminus\{r+1\}$ and $Q_{r+1}$. First, it is guaranteed that $r{+}1\in P_{r+1}$ because of the rule for matching ranks in the critical window from Algorithm 1. Specifically, we ensure that rank r+1r{+}1 is never paired with a rank whose active chunk is $c_{l}$ for $l>r{+}1{-}\log n$. This rule ensures that $r{+}1\in P_{r+1}$ because the only active chunk it could have at the beginning of round r+1r{+}1 is then $c_{r+1-\log n}$, guaranteeing that $r{+}1\in P_{r+1}$.
The next question is whether we can ensure that there exists a matching in round r+1r{+}1. First, we need to ensure that any rank $l$ in the critical window is not matched with a rank whose active chunk will continue to be active by round $l$. If $l\in Q_{r+1}$, this is trivial because the rank does not receive any new active chunk in that round (since it only receives the chunk that is due in that round). If $l\in P_{r+1}$, we just have to ensure that $l$ is not matched with any of the “newest” active chunks; in fact, matching $l$ with ranks possessing the second-oldest active chunk always works to keep $l$ in future $P_{r}$. Since $G_{r}$, the graph that connects $P_{r}$ and $Q_{r}$, is a complete bipartite graph, we can easily guarantee this pairing for any $l$. By removing these $l$ and their partners from their respective sets (either $P_{r}$ or $Q_{r}$), we can trivially apply Hall’s Marriage Theorem since $|P_{r}|=|Q_{r}|$ (and $G_{r}$ is bipartite) to verify that a matching exists. Hence, $\mathcal{I}(r{+}1)$ holds, completing the inductive step. ∎
@@
After round n2n{-}2, the straggler, $\sigma$, now has all fully reduced chunks, and is no longer constrained by whom to pair with. This means we can add the straggler to $Q_{r}$ for subsequent rounds $r\geq n{-}1$ such that $|P_{r}|=|Q_{r}|=\tfrac{n}{2}$. (We enforce the invariant for subsequent rounds that the straggler can only send $c_{n-2}$.) Both ranks n2n{-}2 and $\sigma$ are in $Q_{r}$ (and will remain in $Q_{r}$ until the end of the algorithm), and their active chunk is $c_{n-2}$. Thus, $c_{n-2}$ begins with 2 ranks holding it, as opposed to 1, jumpstarting its doubling process. Every chunk can continue to double because $|P_{r}|=|Q_{r}|$ and every rank only has one active chunk, so via Hall’s Marriage Theorem a perfect matching exists. Therefore, only $\log n-1$ rounds are required to fully propagate the final chunk.
@@
Naturally, n1n{-}1 rounds are required for all chunks to become fully reduced, via the linear sequence of pairing with the straggler, with chunk $c_{j}$ fully reduced in round $j$. Lemma 2 proves that each chunk $c_{j}$ propagates by round $j{+}\log n$. Thus, the total number of rounds in the algorithm is bounded by the final chunks. Chunk $c_{n-3}$ will propagate by round $n{-}3{+}\log n$ and chunk $c_{n-2}$ will have fully propagated after round $n{-}2{+}\log n{-}1=n{-}3{+}\log n$. Since we zero-index rounds, this results in $n{+}\log n{-}2$ rounds in total. ∎
Appendix D StragglAR for Non-Power-of-$2$ Values of $n$
@@
When the number of GPUs is not a power of $2$, StragglAR becomes more challenging to implement because GPUs can possess multiple active chunks at any point in time. To handle the even, non-power-of-$2$ case, we use a similar matching-based approach to construct the schedule in every round, but instead connect vertices $u$ and $v$ with weight $2$ if $u$ possesses a chunk that $v$ needs and $v$ possesses a chunk that $u$ needs, and with weight $1$ if only $u$ possesses a chunk that $v$ needs, but the opposite is not true (or vice versa). Then, we run well-known polynomial-time algorithms [Edmonds, 1965a, b] to compute the maximum weight matchings per round in order to generate the schedule. Thus, offline schedule generation when $n$ is not a power-of-2 has a higher runtime, but is still polytime computable.
@@
For even, non-power-of-2 values of $n$, we do not prove a bound on the $\alpha{-}\beta$ cost. However, we still synthesize and evaluate these schedules in §4 and typically find schedules that complete in ${\sim}n{+}2\log n{-}2$ rounds in practice, which still results in lower $\beta$ cost than baselines. The realized communication complexity of the schedules generated by StragglAR for non-power-of-2 $n$ is shown in the ideal-case performance of Fig. 8, and still outperforms all baselines using the analytical network simulator. We note that StragglAR does not provide schedules for odd values of $n$, but these settings are less common in large-scale distributed ML; further, StragglAR’s schedules for even $n$ are all computable in polytime due to the graph matching structure.
We compare StragglAR’s ideal and worst-case performance to baselines as the cluster size scales, in simulation with the analytical network model. We report the results in Fig. 8, which shows a similar pattern as for powers-of-2. In ideal scenarios, when the ReduceScatter precondition can be fully overlapped with the straggler delay, StragglAR significantly outperforms all baselines. For worst-case scenarios with no straggler delay, StragglAR’s performance is slightly worse than baselines at small cluster sizes, but performs equally well at large cluster sizes due to its $2\times$ lower asymptotic communication complexity.
Figure 8: AllReduce performance of different algorithms as cluster size $n$ scales, for even values of $n$ that are not powers of 2. The shaded region captures the range of StragglAR’s possible performance, ranging from worst case (no straggler and hence no overlap) to ideal (full overlap with straggler).
Appendix E Discussion of Baselines
We provide in-depth explanations and $\alpha{-}\beta$ costs of the baselines. Classical algorithms (e.g., Ring, RH/D) achieve the known lower bounds for AllReduce in networks with homogeneous bandwidth and thus remain the standard implementations in state-of-the-art CCLs [Hu et al., 2025]. While newer works synthesize algorithms for irregular topologies and heterogeneous networks [Won et al., 2024, Wang et al., 2020, Zhao et al., 2024a], they do not provide new algorithms for homogeneous scale-up networks, aside from MSCCL [Cowan et al., 2023], which we compare to.
Ring [Patarasuk and Yuan, 2009]. This is the standard bandwidth-efficient algorithm implemented by NCCL for AllReduce. It divides the buffer into chunks of size $\tfrac{s}{n}$ and consists of 2(n1)2(n-1) rounds in a ring-like communication pattern. The total runtime is $T_{Ring}=2(n-1)\cdot\alpha+2\tfrac{n-1}{n}s\cdot\beta$.
Recursive halving/doubling [RHD] [Bruck et al., 1994]. Also called the Butterfly algorithm, this algorithm—like Ring—consists of two phases. In the first phase, the buffer is first divided in $1/2$ with one partner, then in $1/4$ with another partner, etc. to achieve a ReduceScatter. Then, a mirror-image AllGather completes to propagate all fully reduced chunks. The total runtime is $T_{RHD}=2\log n\cdot\alpha{+}2\tfrac{n-1}{n}s\cdot\beta$.
Broadcast. This is a straggler-aware baseline that assumes the non-stragglers complete an AllReduce ($2\times$ the time as ReduceScatter) during the straggler delay. After the straggler is ready, it exchanges its entire buffer with any other rank to fully reduce the entire buffer and then initiates a broadcast with $\log n$ rounds and $s$ bytes per round. The total ideal-case runtime is $T_{Bcast}=\log n\cdot\alpha+\log n\cdot s\cdot\beta$ while the worst-case runtime is $(2n+\log n-4)\cdot\alpha+(\log n+2\frac{n-2}{n-1})\cdot s\cdot\beta$.
MSCCL [Cowan et al., 2023]. This is a new algorithm synthesized for switched scale-up domains, such as NVIDIA DGX. The algorithm, called allpairs, consists of just two rounds. The buffer is divided into $n$ chunks. In the first round, each GPU is responsible for reducing one chunk (i.e., ReduceScatter), and sends all other chunks to other GPUs simultaneously, thereby splitting the link bandwidth among n1n{-}1 parallel connections. The next step is a mirror-image AllGather, where all GPUs communicate simultaneously to receive all reduced chunks. Thus, the total runtime is $T_{MSCCL}=2\cdot\alpha{+}2\tfrac{n-1}{n}s\cdot\beta$.
Appendix F Evaluation on Perlmutter with 4 GPUs
@@
We show the results from running on one node of the Perlmutter supercomputer [NERSC, 2025] in Fig. 9. While StragglAR still consistently outperforms baselines across buffer sizes in the 4-GPU setup, its gains are less substantial than for the 8-GPU setup. This is expected, and the results are in line with the theoretical results in §4 because StragglAR’s performance advantage improves as $n$, the size of the GPU cluster, scales. (Intuitively, this is because the proportion of “work” that is offloaded to the precondition is higher as $n$ scales.) Fig. 9b indicates similar performance on the 4-GPU node even with both the average and masked delays, suggesting the ReduceScatter precondition can be effectively masked in typical scenarios.
Appendix G NCCL Tuning Behavior
As we discuss in §4, the performance of ncclSend()/ncclRecv() deviates from the $\alpha{-}\beta$ cost model’s expected performance for some specific buffer sizes due to heuristics and protocol-switching behavior implemented internally by NCCL. We run the baseline nccl-tests [NVIDIA, 2024b], a benchmarking suite provided by NVIDIA for evaluating in-built NCCL performance, across different data sizes and plot the results in Fig. 10. The region from 64 MiB-256 MiB deviates from expected behavior under $\alpha{-}\beta$ costs because sending $2\times$, and even $4\times$, the amount of data results in almost no change in communication time with the P2P API. (Under the $\alpha{-}\beta$ model, the time should consistently scale linearly with the send/recv data size, instead of step-wise as shown in Fig. 10.) This is why the Direct and MSCCL AllReduce baselines outperform other algorithms at 256 MiB, as the time to send an entire 256 MiB buffer is only marginally higher than sending smaller chunk sizes, which bandwidth-efficient algorithms like Ring, RHD, and StragglAR do. We do not observe this finding to generalize to other architectures, and it may be a relic of the NCCL version, CUDA version, and specific environment configured by RunPod for ncclSend()/ncclRecv(). Also, outside of the 64-256 MiB range, the performance closely reflects the $\alpha{-}\beta$ model, and our findings in §4 reflect this. This anomalous behavior is also well documented by recent works, which even propose methods to optimize tuning parameters for NCCL [Hu et al., 2025, Wang et al., 2025].
@@ Appendix G NCCL Tuning Behavior
Appendix H End-to-End Training Experiments on Multi-GPU Servers
@@
We evaluate StragglAR using data-parallel (DP) fine-tuning of the Llama3.2-3B model in PyTorch on a DGX A100 VM (8-GPU setup) on RunPod. In this evaluation, we enable AllReduce algorithms to run in the wild, calling implementations directly from PyTorch on a real hardware testbed. To do so, we compile our C++ and CUDA implementations (using the NCCL P2P API, exactly as described in §4) into a custom package, and enable dist.all_reduce() to call these implementations using PyTorch Distributed’s C++ extension capabilities [PyTorch, 2022]. We compare StragglAR to the Ring algorithm, the most well-known and widely used bandwidth-efficient baseline, under this scheme. Using standard PyTorch profiling tools, we identify device (GPU) 6 to be the most frequent straggler on our server, as device 6 was most frequently the straggler in the profiling logs. At the start of training, we pass the straggler rank to the CUDA runtime to ensure StragglAR executes correctly. Then, training simply proceeds as per usual with PyTorch.
@@
Next, we fine-tune Llama3.2-3B using the same data-parallel setup on 8 GPUs with the global batch size of 32 for 100 iterations to determine the overall training efficiency gains with StragglAR. Fig. 11 shows that even these first 100 iterations, StragglAR reduces training time by 7 seconds. These gains suggest that training at longer time scales—for example, days—can shave off several hours of training. Further, we note that StragglAR’s loss curve is just the Ring AllReduce curve shifted to the left: the $y$-coordinates are exactly the same because both algorithms ensure complete data buffers are reduced, ensuring no change in model convergence. Instead, StragglAR enables an unmodified, but faster AllReduce to the application layer, resulting in greater training efficiency and faster time-to-convergence. Fig. 11 captures the wall-clock time, i.e., all compute and communication operations in training, so the results may improve in larger-scale settings (as shown in simulations in §I) or with even larger model sizes where communication becomes a greater bottleneck.
Figure 11: End-to-end evaluation for Llama-3.2-3B fine-tuning on a RunPod DGX A100.
Appendix I End-to-End Training Simulations at Scale